# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### v1

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
$+\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\lVert\nabla u_{N}\rVert_{L^{2}}^{2}+\|u_{N}\|_{L^{2}}^{2}+1,$ (3.17)
where we have used (3.15), the Brézis-Gallouët-Wainger inequality (2.2), as well as Young’s inequality. Note that we have introduced a constant $R_{1}>0$ in order to make the quantity $R_{1}\sigma_{N}$ dimensionless (so that we can apply the Brézis-Gallouët-Wainger inequality). One can choose $R_{1}=\lVert\sigma_{0}\rVert_{L^{\infty}}^{-1}$ for example (for nonzero initial data).
Thanks to the $L^{2}$-estimate in (3.8), we can apply the logarithmic Grönwall inequality from Lemma 2.2 to (3.17) which leads to the following triple-exponential bound
$\sup_{t\in[0,T]}\biggl{[}\|\nabla u_{N}\|_{L^{2}}^{2}+\mathcal{E}^{-1}\|\nabla\sigma_{N}\|_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\|\Delta\sigma_{N}\|_{L^{2}}^{2}\biggr{]}\leq C\exp(\exp(\exp(CT))),$ (3.18)
for some $C$ independent of $N,\epsilon$, and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. Note that the triple-exponential nature of the bound comes from the application of the logarithmic Grönwall inequality to the single-exponential $L^{2}$-estimate (3.8).
$H^{2}$-estimate. One can compute directly that
$\nabla\mathcal{D}_{\epsilon}^{N}$ $=\frac{D(u_{N}):\big{[}\nabla D(u_{N})\big{]}}{\mathcal{D}_{\epsilon}^{N}}=\frac{\frac{1}{2}\nabla\lvert D(u_{N})\rvert^{2}}{\mathcal{D}_{\epsilon}^{N}},$ (3.19)
$\lVert\nabla\mathcal{D}_{\epsilon}^{N}\rVert_{L^{2}}$ $\lesssim\lVert\Delta u_{N}\rVert_{L^{2}},\qquad\text{where the constant is independent of $\epsilon$ and $N$}.$
In addition, we recall the following computation from [53,  p. 22-23], for $i,j=1,2$, with $V_{w}$ given in (3.9),
$\partial^{2}_{ij}\big{(}\lvert V_{w}\rvert V_{w}\big{)}=$ $\;\lvert V_{w}\rvert\partial^{2}_{ij}V_{w}+\frac{V_{w}\cdot\partial_{i}V_{w}}{\lvert V_{w}\rvert}\partial_{j}V_{w}+\frac{V_{w}\cdot\partial_{j}V_{w}}{\lvert V_{w}\rvert}\partial_{i}V_{w}$ (3.20)
@@
$\lesssim\big{(}\lVert V_{w}\rVert_{L^{\infty}}+\lVert V_{w}\rVert_{H^{1}}\big{)}\lVert V_{w}\rVert_{H^{2}},$ (3.21)
where we have applied the Ladyzhenskaya inequality (2.3).
Then, after taking the $L^{2}$-inner product of (3.1a) and (3.1d) with $\Delta^{2}u_{N}$ and $\Delta^{2}\sigma_{N}$, respectively, using integration by parts, and summing up the results, one obtains
$\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\Delta u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\nabla\Delta\sigma_{N}\rVert_{L^{2}}^{2}\bigg{]}$
@@
$+\int_{\mathbb{T}^{2}}\bigg{[}\Delta u_{N}\cdot(\nabla\cdot\Delta\sigma_{N})+\Delta\sigma_{N}:D(\Delta u_{N})\bigg{]}dx$
[…]
@@
$+\|\sigma_{N}\|_{H^{2}}+1\bigg{]}\cdot\big{[}\lVert u_{N}\rVert_{H^{2}}^{2}+\lVert\sigma_{N}\rVert_{H^{3}}^{2}\big{]}+1,$ (3.22)
where we have used (3.21), the Brézis-Gallouët-Wainger inequality (2.2), as well as Young’s inequality. Note that as before, we have used parameters $R_{1},R_{2}>0$ in order to make the quantities $R_{1}\sigma_{N}$ and $R_{2}u_{N}$ dimensionless, so that we can apply the Brézis-Gallouët-Wainger inequality. One can choose $R_{2}=\lVert u_{0}\rVert_{L^{\infty}}^{-1}$ for example (in the case of nonzero initial data).
From (3.2), by using the $L^{2}$-estimate (3.8), the $H^{1}$-estimate (3.18), and the logarithmic Grönwall inequality in Lemma 2.2, we find the following bound:
$\sup_{t\in[0,T]}\biggl{[}\lVert u_{N}(t)\rVert_{H^{2}}^{2}+\lVert\sigma_{N}(t)\rVert_{H^{3}}^{2}\biggr{]}\leq C\exp(\exp(\exp(\exp(\exp(CT))))),$ (3.23)
for some constant $C$ which is independent of $N,\epsilon$, and $T$, but which depends on $\mathcal{E},\alpha$, and the initial data. Note that the quintuple-exponential growth in time of the bound originates from the single-exponential respectively triple-exponential growth in time of the $L^{2}$ and $H^{1}$-estimates, and the subsequent application of the logarithmic Grönwall inequality.
Estimate on the time derivative and summary. Thanks to (3.23) and system (3.1), one can obtain directly that
$\sup_{t\in[0,T]}\biggl{[}\|\partial_{t}u_{N}(t)\|_{H^{2}}^{2}+\|\partial_{t}\sigma_{N}-\alpha^{2}\partial_{t}\Delta\sigma_{N}\|_{H^{1}}^{2}\biggr{]}\leq C\exp(\exp(\exp(\exp(\exp(CT))))).$ (3.24)
Meanwhile, by directly applying integration by parts we deduce that
@@
Hence, equations (3.23) and (3.24) yield
$\begin{gathered}\sup_{t\in[0,T]}\biggl{[}\lVert u_{N}(t)\rVert_{H^{2}}+\lVert\sigma_{N}(t)\rVert_{H^{3}}+\|\partial_{t}u_{N}(t)\|_{H^{2}}+\|\partial_{t}\sigma_{N}\|_{H^{3}}\biggr{]}\\ \leq C\exp(\exp(\exp(\exp(\exp(CT))))),\end{gathered}$ (3.25)
@@
3.3 Well-posedness of the intermediate system (1.13)
Passing the limit $N\rightarrow\infty$ and existence of the strong solution. For any fixed $T\in(0,\infty)$ and $\epsilon\in(0,\infty)$, from the bound (3.25) and by applying the Aubin-Lions compactness lemma (see [9, Theorem II.5.16]), one has the following convergence results: for a suitable subsequence of $N\rightarrow\infty$,
$u_{N}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~{}L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (3.26)
$\partial_{t}u_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~{}L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (3.27)
$u_{N}$ $\rightarrow u$ strongly in $~{}C([0,T];H^{1}(\mathbb{T}^{2})),$ (3.28)
$\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~{}L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (3.29)
$\sigma_{N}$ $\rightarrow\sigma$ strongly in $~{}C([0,T];H^{2}(\mathbb{T}^{2})),$ (3.30)
$\partial_{t}\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~{}L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (3.31)
where $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ with the same estimates as in (3.25) with $(u_{N},\sigma_{N})$ replaced by $(u,\sigma)$, i.e.,
$\begin{gathered}\sup_{t\in[0,T]}\biggl{[}\lVert u(t)\rVert_{H^{2}}+\lVert\sigma(t)\rVert_{H^{3}}+\|\partial_{t}u(t)\|_{H^{2}}+\|\partial_{t}\sigma\|_{H^{3}}\biggr{]}\\ \leq C\exp(\exp(\exp(\exp(\exp(CT))))),\end{gathered}$ (3.32)
for some constant $C$ independent of $\epsilon$ and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. We also have $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ thanks to [66, Lemma 1.3.2, p. 196].
It is clear that one can pass to the limit in the linear terms of the Galerkin system (3.1). To show the convergence of the nonlinear terms, first we observe that $\mathcal{D}_{\epsilon}^{N}\rightarrow\mathcal{D}_{\epsilon}$ in $C([0,T];L^{2}(\mathbb{T}^{2}))$. In fact, due to Lemma 2.3, we have by using (3.28)
@@
Notice that this bound holds pointwise in time, as will generally be the case for the estimates in this part of the paper. We will therefore omit the temporal part of the norm in this section. Next, by using (3.30) one can estimate that
@@
by the convergence of both $\mathcal{D}_{\epsilon}^{N}$ as well as $\sigma_{N}$. Similarly, we know that the other nonlinear terms in the evolution equation for the stress tensor, i.e. $-\frac{3}{2P}\mathcal{D}_{\epsilon}^{N}\operatorname{Tr}\sigma_{N}I$ and $\frac{\mathcal{D}_{\epsilon}^{N}}{2}I$, converge to $-\frac{3}{2P}\mathcal{D}_{\epsilon}\operatorname{Tr}\sigma I$ and $\frac{\mathcal{D}_{\epsilon}}{2}I$ in $L^{\infty}((0,T);L^{2}(\mathbb{T}^{2}))$ respectively.
@@
Therefore all the terms in the Galerkin system (3.1) converge, and $(u,\sigma)$ solves the intermediate system (1.13). This concludes the proof of the existence of a strong solution to the intermediate model (1.13).
@@
Suppose that $(u_{1},\sigma_{1}),(u_{2},\sigma_{2})\in C([0,T];H^{2}(\mathbb{T}^{2}))\times C([0,T];H^{3}(\mathbb{T}^{2}))$ are two strong solutions to the intermediate model (1.13). Let
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (3.33a)
$\frac{1}{\mathcal{E}}\partial_{t}\big{(}\delta\sigma-\alpha^{2}\Delta\delta\sigma\big{)}+\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (3.33b)
$+\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (3.33c)
@@
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big{[}(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big{]},$ (3.33d)
$\mathcal{D}_{\epsilon,i}$ $=\sqrt{\lvert D(u_{i})\rvert^{2}+\epsilon^{2}}.$ (3.33e)
We observe that $\partial_{t}(\delta u)\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$ and $\partial_{t}(\delta\sigma)\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and therefore we can take the $L^{2}$-inner product of equations (3.33a) and (3.33b) with $\delta u$ respectively $\delta\sigma$. We therefore compute the $L^{2}$-estimates on the difference $(\delta u,\delta\sigma)$ as follows (using a cancellation similar to (3.7))
$\begin{gathered}\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg{]}=-\int_{\mathbb{T}^{2}}\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}\operatorname{Tr}\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg{[}\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2})\bigg{]}:\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg{[}\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}\bigg{]}:\delta\sigma\ {\rm d}{x}\\ +\int_{\mathbb{T}^{2}}\big{(}\mathcal{T}_{w,1}-\mathcal{T}_{w,2}\big{)}\cdot\delta u\ {\rm d}{x}=:J_{1}+J_{2}+J_{3}+J_{4}.\end{gathered}$ (3.34)
Applying Hölder’s inequality, inequality (2.7), and the Ladyzhenskaya inequality (2.3) yields
@@
$\begin{gathered}\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg{]}\lesssim(\lVert\sigma_{1}\rVert_{L^{\infty}}+1)\big{[}\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\lVert\delta\sigma\rVert_{L^{2}}^{2}\big{]}\\ +\lVert\nabla u_{2}\rVert_{L^{2}}\big{[}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\big{]}+(\lVert u_{1}\rVert_{L^{\infty}}+\lVert u_{2}\rVert_{L\infty}+\lVert U_{w}\rVert_{L^{\infty}})\lVert\delta u\rVert_{L^{2}}^{2}.\end{gathered}$ (3.35)
Meanwhile, to calculate the $H^{1}$ estimate of the difference $(\delta u,\delta\sigma)$, similar to (3.17), we have (by using a cancellation similar to (3.7))
$\frac{1}{2}\frac{d}{dt}\bigg{[}\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg{]}$ (3.36)
@@
Similarly to the $L^{2}$-estimate, we establish (by using inequality (2.7))

### v2

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
In addition, we recall the following computation from p. 22-23 in Ref. [57], for $i,j=1,2$, with $V_{w}$ given in (33),
$\partial^{2}_{ij}\big(\lvert V_{w}\rvert V_{w}\big)=$ $\;\lvert V_{w}\rvert\partial^{2}_{ij}V_{w}+\frac{V_{w}\cdot\partial_{i}V_{w}}{\lvert V_{w}\rvert}\partial_{j}V_{w}+\frac{V_{w}\cdot\partial_{j}V_{w}}{\lvert V_{w}\rvert}\partial_{i}V_{w}$ (46)
@@
$\lesssim\big(\lVert V_{w}\rVert_{L^{\infty}}+\lVert V_{w}\rVert_{H^{1}}\big)\lVert V_{w}\rVert_{H^{2}},$ (47)
where we have applied the Ladyzhenskaya inequality (17).
Then, after taking the $L^{2}$-inner product of (25a) and (25d) with $\Delta^{2}u_{N}$ and $\Delta^{2}\sigma_{N}$, respectively, using integration by parts, and summing up the results, one obtains
$\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\Delta u_{N}\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\Delta\sigma_{N}\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\nabla\Delta\sigma_{N}\rVert_{L^{2}}^{2}\bigg]$
@@
$+\int_{\mathbb{T}^{2}}\bigg[\Delta u_{N}\cdot(\nabla\cdot\Delta\sigma_{N})+\Delta\sigma_{N}:D(\Delta u_{N})\bigg]\ {\rm d}x$
$\lesssim\,$ $\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\lVert\nabla\sigma_{N}\rVert_{L^{4}}\lVert\Delta\sigma_{N}\rVert_{L^{4}}+\big[\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\lVert\sigma_{N}\rVert_{L^{\infty}}+\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\big]\lVert\nabla\Delta\sigma_{N}\rVert_{L^{2}}$
@@
$+\|\sigma_{N}\|_{H^{2}}+1\bigg]\cdot\big[\lVert u_{N}\rVert_{H^{2}}^{2}+\lVert\sigma_{N}\rVert_{H^{3}}^{2}\big]+1$
$\lesssim$ $\bigg[\lVert\sigma_{N}\rVert_{H^{1}}^{2}+\log\big(e+C_{L}R_{1}\lVert\sigma_{N}\rVert_{H^{2}}\big)+\lVert u_{N}\rVert_{H^{1}}^{2}+\log\big(e+C_{L}R_{2}\lVert u_{N}\rVert_{H^{2}}\big)$
$+\|\sigma_{N}\|_{H^{2}}+1\bigg]\cdot\big[\lVert u_{N}\rVert_{H^{2}}^{2}+\lVert\sigma_{N}\rVert_{H^{3}}^{2}\big]+1,$ (48)
where we have used (47), the Brézis-Gallouët-Wainger inequality (16), the Sobolev embedding theorem as well as Young’s inequality. The terms on the second line are negative definite, and in the seventh line we estimate the term $\lVert\nabla^{2}u_{N}\rVert_{L^{2}}\lVert\nabla\sigma_{N}\rVert_{L^{4}}\lVert\Delta\sigma_{N}\rVert_{L^{4}}$ by means of the Sobolev embedding theorem, which gives $\lVert u_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{2}}\lVert\sigma_{N}\rVert_{H^{3}}$. Applying Young’s inequality leads to some of the terms on the last line of estimate (3.2). Due to the monotonicity of the logarithm, we can replace the arguments $\lVert\sigma_{N}\rVert_{H^{2}}$ and $\lVert u_{N}\rVert_{H^{2}}$ by their sum, which facilitates the application of the logarithmic Grönwall inequality.
Note that as before, we have used parameters $R_{1},R_{2}>0$ in order to make the quantities $R_{1}\sigma_{N}$ and $R_{2}u_{N}$ dimensionless (such that the resulting quantities do not have physical units), so that we can apply the Brézis-Gallouët-Wainger inequality. One can choose $R_{2}=\lVert u_{0}\rVert_{L^{\infty}}^{-1}$ for example (in the case of nonzero initial data). Moreover, similarly to before we have used the following cancellation
$\int_{\mathbb{T}^{2}}\bigg[\Delta u_{N}\cdot(\nabla\cdot\Delta\sigma_{N})+\Delta\sigma_{N}:D(\Delta u_{N})\bigg]\ {\rm d}x=0.$ (49)
From (3.2), by using the $L^{2}$-estimate (32), the $H^{1}$-estimate (44), and the logarithmic Grönwall inequality in Lemma 2.2, we find the following bound:
$\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}^{2}+\lVert\sigma_{N}(t)\rVert_{H^{3}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))),$ (50)
for some constant $C$ which is independent of $N,\epsilon$, and $T$, but which depends on $\mathcal{E},\alpha$, and the initial data. Note that the triple-exponential growth in time of the bound originates from the single-exponential respectively double-exponential growth in time of the $L^{2}$ and $H^{1}$-estimates, and the subsequent application of the logarithmic Grönwall inequality.
Estimate on the time derivative and summary. Thanks to (50) and system (25), one can obtain directly that
$\sup_{t\in[0,T]}\biggl[\|\partial_{t}u_{N}(t)\|_{H^{2}}^{2}+\|\partial_{t}\sigma_{N}-\alpha^{2}\partial_{t}\Delta\sigma_{N}\|_{H^{1}}^{2}\biggr]\leq C\exp(\exp(\exp(CT))).$ (51)
This estimate follows because all the other terms (i.e. except the time derivative) in Eq. (25a) can be shown to lie in $C([0,T];H^{2}(\mathbb{T}^{2}))$, while all the other terms in Eq. (25d) are elements of $C([0,T];H^{1}(\mathbb{T}^{2}))$ (because of estimate (50)). Meanwhile, by directly applying integration by parts we deduce that
@@
Hence, equations (50) and (51) yield
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u_{N}(t)\rVert_{H^{2}}+\lVert\sigma_{N}(t)\rVert_{H^{3}}+\|\partial_{t}u_{N}(t)\|_{H^{2}}+\|\partial_{t}\sigma_{N}\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (52)
@@
3.3 Well-posedness of the intermediate system (13)
Passing the limit $N\rightarrow\infty$ and existence of the strong solution. For any fixed $T\in(0,\infty)$ and $\epsilon\in(0,\infty)$, from the bound (52) and by applying the Aubin-Lions compactness lemma (see Theorem II.5.16 in Ref. [9]), one has the following convergence results: for a suitable subsequence of $N\rightarrow\infty$,
$u_{N}$ $\overset{\ast}{\rightharpoonup}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (53)
$\partial_{t}u_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}u$ weakly-* in $~L^{\infty}((0,T);H^{2}(\mathbb{T}^{2})),$ (54)
$u_{N}$ $\rightarrow u$ strongly in $~C([0,T];H^{1}(\mathbb{T}^{2})),$ (55)
$\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (56)
$\sigma_{N}$ $\rightarrow\sigma$ strongly in $~C([0,T];H^{2}(\mathbb{T}^{2})),$ (57)
$\partial_{t}\sigma_{N}$ $\overset{\ast}{\rightharpoonup}\partial_{t}\sigma$ weakly-* in $~L^{\infty}((0,T);H^{3}(\mathbb{T}^{2})),$ (58)
where $u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\partial_{t}u\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$, $\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and $\partial_{t}\sigma\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$ with the same estimates as in (52) with $(u_{N},\sigma_{N})$ replaced by $(u,\sigma)$, i.e.,
$\begin{gathered}\sup_{t\in[0,T]}\biggl[\lVert u(t)\rVert_{H^{2}}+\lVert\sigma(t)\rVert_{H^{3}}+\|\partial_{t}u(t)\|_{H^{2}}+\|\partial_{t}\sigma\|_{H^{3}}\biggr]\\ \leq C\exp(\exp(\exp(CT))),\end{gathered}$ (59)
for some constant $C$ independent of $\epsilon$ and $T$, but depending on $\mathcal{E},\alpha$, and the initial data. We also have $u\in C([0,T];H^{2}(\mathbb{T}^{2}))$ and $\sigma\in C([0,T];H^{3}(\mathbb{T}^{2}))$ thanks to Lemma 1.3.2 on p. 196 in Ref. [70].
It is clear that one can pass to the limit in the linear terms of the Galerkin system (25). To show the convergence of the nonlinear terms, first we observe that $\mathcal{D}_{\epsilon}^{N}\rightarrow\mathcal{D}_{\epsilon}$ in $C([0,T];L^{2}(\mathbb{T}^{2}))$. In fact, due to Lemma 2.3, we have by using (55)
@@ 3.3 Well-posedness of the intermediate system (13)
Notice that this bound holds pointwise in time, as will generally be the case for the estimates in this part of the paper. We will therefore omit the temporal part of the norm in this section. Next, by using (57) one can estimate that
@@
by the convergence of both $\mathcal{D}_{\epsilon}^{N}$ as well as $\sigma_{N}$. Similarly, we know that the other nonlinear terms in the evolution equation for the stress tensor, i.e. $\frac{3}{2P}\mathcal{D}_{\epsilon}^{N}\operatorname{Tr}\sigma_{N}I$ and $\frac{\mathcal{D}_{\epsilon}^{N}}{2}I$, converge to $\frac{3}{2P}\mathcal{D}_{\epsilon}\operatorname{Tr}\sigma I$ and $\frac{\mathcal{D}_{\epsilon}}{2}I$ in $L^{\infty}((0,T);L^{2}(\mathbb{T}^{2}))$ respectively.
@@
Therefore all the terms in the Galerkin system (25) converge, and $(u,\sigma)$ solves the intermediate system (13). This concludes the proof of the existence of a strong solution to the intermediate model (13).
@@
Suppose that $(u_{1},\sigma_{1}),(u_{2},\sigma_{2})\in C([0,T];H^{2}(\mathbb{T}^{2}))\times C([0,T];H^{3}(\mathbb{T}^{2}))$ are two strong solutions to the intermediate model (13). Let
@@
$\partial_{t}\delta u=\nabla\cdot\delta\sigma+\mathcal{T}_{w,1}-\mathcal{T}_{w,2}+\Omega(\delta u)^{\perp},$ (60a)
$\frac{1}{\mathcal{E}}\partial_{t}\big(\delta\sigma-\alpha^{2}\Delta\delta\sigma\big)+\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}$ $\mathbb{I}_{2})$ (60b)
$+\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}+\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}$ $\mathbb{I}_{2}=D(\delta u),$ (60c)
@@
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (60d)
$\mathcal{D}_{\epsilon,i}$ $=\sqrt{\lvert D(u_{i})\rvert^{2}+\epsilon^{2}}.$ (60e)
We observe that $\partial_{t}(\delta u)\in L^{\infty}((0,T);H^{2}(\mathbb{T}^{2}))$ and $\partial_{t}(\delta\sigma)\in L^{\infty}((0,T);H^{3}(\mathbb{T}^{2}))$, and therefore we can take the $L^{2}$-inner product of equations (60a) and (60b) with $\delta u$ respectively $\delta\sigma$. We therefore compute the $L^{2}$-estimates on the difference $(\delta u,\delta\sigma)$ as follows (using a cancellation similar to (31))
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]=-\int_{\mathbb{T}^{2}}\frac{\mathcal{D}_{\epsilon,1}-\mathcal{D}_{\epsilon,2}}{2}\operatorname{Tr}\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{4\mathcal{D}_{\epsilon,1}}{P}(\sigma_{1}-\frac{1}{2}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2})-\frac{4\mathcal{D}_{\epsilon,2}}{P}(\sigma_{2}-\frac{1}{2}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2})\bigg]:\delta\sigma\ {\rm d}{x}\\ -\int_{\mathbb{T}^{2}}\bigg[\frac{\mathcal{D}_{\epsilon,1}}{2P}\operatorname{Tr}\sigma_{1}\mathbb{I}_{2}-\frac{\mathcal{D}_{\epsilon,2}}{2P}\operatorname{Tr}\sigma_{2}\mathbb{I}_{2}\bigg]:\delta\sigma\ {\rm d}{x}\\ +\int_{\mathbb{T}^{2}}\big(\mathcal{T}_{w,1}-\mathcal{T}_{w,2}\big)\cdot\delta u\ {\rm d}{x}=:J_{1}+J_{2}+J_{3}+J_{4}.\end{gathered}$ (61)
Applying Hölder’s inequality, inequality (21), and the Sobolev embedding theorem yields
@@
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\bigg]\lesssim(\lVert\sigma_{1}\rVert_{L^{\infty}}+1)\big[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\lVert\delta\sigma\rVert_{L^{2}}^{2}\big]\\ +\lVert\nabla u_{2}\rVert_{L^{2}}\big[\lVert\delta\sigma\rVert_{L^{2}}^{2}+\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}\big]+(\lVert u_{1}\rVert_{L^{\infty}}+\lVert u_{2}\rVert_{L\infty}+\lVert U_{w}\rVert_{L^{\infty}})\lVert\delta u\rVert_{L^{2}}^{2}.\end{gathered}$ (62)
Meanwhile, to calculate the $H^{1}$ estimate of the difference $(\delta u,\delta\sigma)$, similar to (42), we have (by using a cancellation similar to (31))
$\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\alpha^{2}\mathcal{E}^{-1}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]$ (63)
@@
Similarly to the $L^{2}$-estimate, we establish (by using inequality (21) and the Sobolev embedding theorem)